OpenStudy (anonymous):

solve for quadratic equ 2x^2 - 4x + 1 = 0

OpenStudy (anonymous):

oops nvm

OpenStudy (anonymous):

Do you know what a= b= and c=

OpenStudy (anonymous):

am going to explain to you detailed every exercise of the following way: First exercise: 2x² - 4x + 1 = 0 I am going to apply the general formula where the values are: a=2 b =-4 c=1 Chord to the quadratic form ax² + bx+c. The formula is: ....-b ± √b²-4ac x=------------------ ............2a Now we replace the values, this way: ....-(-4) ± √(-4)²-4(2)(1) x=---------------------------- ..................2(4) We solve what is in of the square, like that root: .....4 ± √16-8 x=---------------- ...........8 .....4 ± √8 x=------------ ..........8 We rationalize, though really it does not have square exact root and it would remain like that: ....4 ± 2√2 x=------------ ..........8 Now we look for both values for both signs independently, of the following way: And the looked results are: x1 = 4 + 2√2 / 8 x2= 4 - 2√2 / 8 Second exercise: 3x² + 2x - 7 = 0 I am going to apply the general formula where the values are: a=3 b =2 c=-7 Chord to the quadratic form ax² + bx+c. The formula is: ....-b ± √b²-4ac x=------------------ ............2a Now we replace the values, this way: ....-(2) ± √(2)²-4(3)(-7) x=---------------------------- ................2(-7) We solve what is in of the square, like that root: .....-2 ± √4+84 x=---------------- ...........-14 .....-2 ± √88 x=------------ ........-14 We rationalize, though really it does not have square exact root and it would remain like that: ....-2 ± 2√22 x=-------------- ..........-14 Now we look for both values for both signs independently, of the following way: And the looked results are: x1 = -2 + 2√22 / -14 = 2 - 2√22 /14 x2= -2 - 2√22 / -14 = 2 + 2√22 / 14 I hope that my explanations you are of usefulness! Greet from Colombia, on happy Friday and bye-bye =). Source(s): University student of Civil Engineering I walk of vacations still Experience pesonal and without web

OpenStudy (anonymous):

quadratic formula for this one

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

i just need to have it factored but i think its prime

OpenStudy (anonymous):

you cannot factor using integers, so you need to use the quadratic formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \(a=2,b=-4,c=1\)

OpenStudy (anonymous):

follow @linda18, she has everything I was to lazy to write lol

OpenStudy (anonymous):

i amm

OpenStudy (anonymous):

make the number replacements and get \[x=\frac{4\pm\sqrt{(-4)^2-4\times 2\times 4}}{2\times 1}\] then some careful arithmetic \[x=\frac{4\pm\sqrt{16-8}}{4}=\frac{4\pm\sqrt{8}}{4}\] \[x=\frac{4\pm2\sqrt{2}}{4}\] \[x=\frac{2(2\pm\sqrt{2})}{4}=\frac{2\pm\sqrt{2}}{2}\]

OpenStudy (anonymous):

so x is\[\sqrt{2}\]

OpenStudy (anonymous):

use this formula : -b ± √b²-4ac x=------------------ 2a it will help u

OpenStudy (anonymous):

yeah pm 1.41

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

welcum =)